### v1

# Quantum Origin of Classical Background Fields from Coherent States

Abstract
Classical background electromagnetic fields are routinely employed in quantum electrodynamics to describe a wide range of physical situations, from laser-matter interactions to strong-field phenomena. In this work, we present a first-principles formulation that clarifies the quantum origin of such classical background fields in QED by systematically deriving them from coherent states of the electromagnetic field.
Starting from the operator formulation of QED, we show how scattering amplitudes between coherent states naturally lead to an effective description in terms of background fields, while maintaining a clear separation between the coherent laser mode and other quantized photon degrees of freedom. This framework allows one to consistently incorporate effects beyond the fixed background approximation, such as depletion and backreaction, without assuming any particular field strength or intensity regime.
We further demonstrate how the conventional generating functional with a prescribed background field emerges as a limiting case, corresponding to fixed coherent state boundary conditions. The path integral representation is then obtained as a reformulation of the same underlying Heisenberg picture amplitudes, providing a unified view of operator-based and functional approaches.
Our results establish a general and intensity-independent foundation for QED with coherent background fields, within which the standard formulations of strong-field QED arise as well-defined special cases.
322-4 Oroshi-cho, Toki, Gifu, 509-5292, Japan
December 24, 2025
1 Introduction
The use of classical background electromagnetic (EM) fields is a standard and highly successful strategy in quantum electrodynamics (QED). It underlies a wide range of applications, including laser-matter interactions [Salamin2006, Joachain_Kylstra_Potvliege_2011, Raicher2014], external-field problems [Greiner_Muller_Rafelski_1985], and strong-field QED (also known as nonlinear QED) [Ritus1985, Piazza2012, Fedotov2023]. In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations are quantized on top of it. Despite its practical success, this raises a conceptual question: how does a classical background field emerge from the underlying quantum theory of the EM field, and under what conditions is the fixed background approximation justified?
In conventional treatments, this issue is often bypassed by postulating the background field from the outset, leading to the framework of QED in external fields, or the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. Within this framework, scattering amplitudes are computed in the presence of a given background, and techniques such as Volkov solutions [Joachain_Kylstra_Potvliege_2011, Ritus1985, Piazza2012, Fedotov2023, Volkov1936, Itzykson-Zuber1980] and background-dependent generating functionals [Greiner_Muller_Rafelski_1985, Schwinger1951] are employed. However, the background field itself is not dynamical, and its relation to the quantized EM field remains implicit.
A natural candidate for bridging this gap is provided by coherent states of the EM field [Glauber1963a, Glauber1963b, Sudarshan1963]. Coherent states exhibit classical-like properties while remaining fully quantum mechanical, and they have long been recognized as an appropriate description of ideal laser fields [Sargent1974, Loudon_2000]. From this perspective, a classical background field may be viewed not as an external input but as an effective description arising from scattering of coherent states of the quantized EM field, see Ref. [Frantz1965] as an earlier work in that research direction.
In this paper, we develop a first-principles formulation of QED in which classical background fields are systematically derived from coherent states, without assuming any specific field strength or intensity regime. In what follows, we refer to this formulation as QED with coherent background fields (hereafter referred to as coherent-field QED), to emphasize that the classical backgrounds arise from coherent states of the quantized EM field rather than being imposed externally. Our approach starts from the operator formulation of QED in the Schrödinger and Heisenberg pictures [Itzykson-Zuber1980, Sakurai1967], and considers scattering amplitudes between asymptotic coherent states of the EM field, together with arbitrary in- and out-states of other particles. This construction allows us to maintain a clear separation between the coherent laser mode and other quantized photon degrees of freedom, and to consistently discuss effects beyond the fixed background approximation, such as depletion and backreaction [Ilderton-Seipt2018]. We show that the conventional generating functional of QED with a prescribed background field [Greiner_Muller_Rafelski_1985, Schwinger1951] emerges as a limiting case, corresponding to fixed coherent state boundary conditions. Furthermore, we reformulate the same scattering amplitudes in terms of a path integral representation, demonstrating that the operator-based and functional approaches describe the same physics from complementary perspectives.
While our formulation is not restricted to the high-intensity regime [Ritus1985, Piazza2012, Fedotov2023], it naturally reproduces standard strong-field QED as well-defined limiting cases. In this sense, the present work provides a unified and intensity-independent foundation for QED in background fields, clarifying the conceptual status of classical backgrounds and their domain of validity. The present formulation provides a first-principles reinterpretation of the fixed background approximation commonly employed in strong-field QED (i.e., within the coherent-field QED framework in the fixed-background limit), by relating it explicitly to coherent quantum states of the EM field.
This paper is organized as follows. In Sec. 2, we introduce coherent states of EM fields. Section 3 presents the formalism connecting the Schrödinger and Heisenberg pictures. Section 4 discusses the Furry picture and Volkov solutions, while Sec. 5 applies the formalism to the Heisenberg–Euler vacuum and laser depletion effects. Section 6 relates the approach to Becchi–Rouet–Stora–Tyutin (BRST) quantization, and Sec. 7 presents the path integral formulation. We conclude with a summary and outlook in Sec. 8.
2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition [Gupta1950, Bleuler1950]. Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
$\partial_{\mu}A_{(+)\mathrm{H}}^{\mu}(x)|\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (1)
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}A_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adapt the Gupta–Bleuler condition to a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state [Glauber1963a, Glauber1963b, Sudarshan1963, Sargent1974, Loudon_2000] is specified by the following eigen equation
$\hat{A}_{(+)\mathrm{H}}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle$ $=\mathcal{A}_{(+)}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle,$ (2)
and that conjugate equation is $\langle\alpha_{\mathrm{H}}|\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=\langle\alpha_{\mathrm{H}}|\mathcal{A}_{(-)}^{\mu}(x)$ for $\mathcal{A}_{(-)}^{\mu}(x)=[\mathcal{A}_{(+)}^{\mu}(x)]^{*}$ which is the complex conjugate of $\mathcal{A}_{(+)}^{\mu}(x)$. Hence, the expectation value of the EM field in a coherent state is
$\mathcal{A}^{\mu}(x)$ $=\langle\alpha_{\mathrm{H}}|\hat{A}_{\mathrm{H}}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle$ (3)
$=\mathcal{A}_{(+)}^{\mu}(x)+\mathcal{A}_{(-)}^{\mu}(x),$ (4)
which, by identifying $|\mathrm{Physical}_{\mathrm{H}}\rangle=|\alpha_{\mathrm{H}}\rangle$ in the Gupta–Bleuler condition (1), automatically satisfies the Lorenz gauge condition
$\partial_{\mu}\mathcal{A}^{\mu}(x)$ $=0.$ (5)
The coherent state can be generated from the vacuum state $|0_{\mathrm{H}}\rangle$ using the displacement operator $\hat{D}_{\mathrm{H}}(\alpha,t)$, so that
$|\alpha_{\mathrm{H}}\rangle$ $=\hat{D}_{\mathrm{H}}(\alpha,t)|0_{\mathrm{H}}\rangle.$ (6)
Applying the displacement operator to the EM field operator yields
$\hat{D}_{\mathrm{H}}^{\dagger}(\alpha,t)\hat{A}_{\mathrm{H}}^{\mu}(x)\hat{D}_{\mathrm{H}}(t,\alpha)$ $=\hat{A}_{\mathrm{H}}^{\mu}(x)+\mathcal{A}^{\mu}(x)\hat{\mathbb{I}}$, (7)
where, $\hat{\mathbb{I}}$ is the identity operator [Glauber1963b, Loudon_2000, Frantz1965]. In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
3 Basic formalism
In this section, we clarify how a classical background field is consistently incorporated within the present operator-based formulation, with particular emphasis on the role of the Schrödinger and Heisenberg pictures. The purpose of this section is twofold. First, we demonstrate how a coherent background field can be introduced in the Schrödinger picture through a displacement operator acting on the quantum gauge field, leading to a clean separation between classical and quantum degrees of freedom at the operator level. Second, we show how the time-dependent background field employed in coherent-field QED emerges naturally upon transforming to the Heisenberg picture. This analysis makes explicit that the time dependence of background fields is a picture-dependent notion rather than a fundamental property of the underlying Hamiltonian. This picture change makes explicit whether the time dependence of the background field is carried by the state or by the operators.
In relevant experiments at high-intensity laser facilities [ELI-NP_2018, CoReLS_Compton_2024, Sarri2025], scattering processes in strong-field QED are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background [Greiner_Muller_Rafelski_1985, Ritus1985, Piazza2012, Moortgat-Pick2009, Hartin2018]. It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian.
From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. As a result, the fixed background approximation, though highly effective in practice, appears conceptually ad hoc when viewed from first principles of quantum field theory.
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states [Itzykson-Zuber1980, Sakurai1967]. Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
Although the Hamiltonian governing the full quantum dynamics is time independent as an operator, the explicit time dependence of the background field commonly used in coherent-field QED arises upon transforming to the Heisenberg picture. In this picture, the time evolution of operators induces a spacetime dependence in the expectation value of the gauge field taken in a coherent state. Thus, the classical background field appearing in the Furry picture should be understood as an emergent, picture-dependent object rather than as a fundamental ingredient of the theory.
This observation provides a resolution of the apparent tension between the use of time-dependent background fields and the time-independent nature of the underlying quantum Hamiltonian. It also clarifies the precise sense in which coherent states serve as the bridge between fully quantized gauge fields and their classical counterparts in coherent-field QED. In the following, we make these statements explicit by constructing the displaced Hamiltonian in the Schrödinger picture and then translating the resulting expressions into the Heisenberg picture, where the connection to the standard background-field formulation becomes manifest.
3.1 Departure from the Schrödinger picture
We consider QED in the presence of a background laser EM field, aiming to reformulate it into coherent-field QED. To this end, we assume that a coherent background laser field is always present. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
$|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{D}_{\mathrm{S}}(\alpha)|\Psi_{\mathrm{S}}(t)\rangle,$ (8)
corresponding to the background field approximation.
The “usual” QED Hamiltonian and its density for the state $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ in the Schrödinger picture are expressed abstractly as
$\hat{H}_{\mathrm{S}}$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$ (9)
$=\int d^{3}\boldsymbol{x}\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{S}}(\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{S}}(\boldsymbol{x}),\hat{A}_{\mathrm{S}}(\boldsymbol{x})),$ (10)
$\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ $=\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{S}}(\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{S}}(\boldsymbol{x}),\hat{A}_{\mathrm{S}}(\boldsymbol{x})).$ (11)
Both $\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ and $\hat{H}_{\mathrm{S}}=\int d^{3}\boldsymbol{x}\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ are taken to be time-independent in the Schrödinger picture [Itzykson-Zuber1980, Sakurai1967]. Under this preparation, the Schrödinger equation governing the “coherent laser field plus other particles” state is
$i\hbar\partial_{t}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{H}_{\mathrm{S}}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle.$ (12)
Formally,
$|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{usual\,QED}}(t,t_{0})|\alpha_{\mathrm{S}}(t_{0}),\Psi_{\mathrm{S}}(t_{0})\rangle,$ (13)
$\hat{U}_{\mathrm{usual\,QED}}(t,t_{0})$ $=\exp\bigg[-\frac{i}{\hbar}\hat{H}_{\mathrm{S}}(t-t_{0})\bigg].$ (14)
We now separate $|\alpha_{\mathrm{S}}(t)\rangle$ and focus on the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$. Substituting Eq. (8) into Eq. (12), the evolution equation becomes
$i\hbar\partial_{t}|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{H}_{\mathrm{S}}\hat{D}_{\mathrm{S}}(\alpha)|\Psi_{\mathrm{S}}(t)\rangle$
$=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}\hat{D}_{\mathrm{S}}(\alpha)]|\Psi_{\mathrm{S}}(t)\rangle.$ (15)
Using the formula (7) in the present picture,
$\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$ $=\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})+\mathcal{A}(t_{0},\boldsymbol{x})\hat{\mathbb{I}}$, (16)
$\mathcal{A}(t_{0},\boldsymbol{x})$ $=\langle\alpha_{\mathrm{S}}(t_{0})|\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})|\alpha_{\mathrm{S}}(t_{0})\rangle,$ (17)
we conclude
$i\hbar\partial_{t}|\Psi_{\mathrm{S}}(t)\rangle$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]|\Psi_{\mathrm{S}}(t)\rangle.$ (18)
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) [Gupta1950, Bleuler1950] evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture, the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
$|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{S}}(t,t_{0})|\Psi_{\mathrm{S}}(t_{0})\rangle.$ (19)
The associated evolution operator satisfies
$i\hbar\partial_{t}\hat{U}_{\mathrm{S}}(t,t_{0})$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0}),$ (20)
with that formal solution
$\hat{U}_{\mathrm{S}}(t,t_{0})$ $=\exp\bigg[-\frac{i}{\hbar}H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}](t-t_{0})\bigg].$ (21)
Thus the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$, with the coherent laser field factored out, incorporates the influence of the laser ($\mathcal{A}$) through its explicit appearance in the evolution operator.
This confirms that the Schrödinger picture evolution operator consistently incorporates the coherent background field through its displacement, while maintaining full quantum dynamics for all other degrees of freedom. In this way, the framework provides a rigorous starting point for connecting the Schrödinger picture with the Heisenberg picture in the presence of a coherent laser background.
3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As declared earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified [Itzykson-Zuber1980, Sakurai1967]. Thus
$|\Psi_{\mathrm{H}}\rangle$ $=|\Psi_{\mathrm{S}}(t_{0})\rangle,$ (22)
and for any observable $\hat{O}$,
$\hat{O}_{\mathrm{H}}(t_{0},\boldsymbol{x})$ $=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x}).$ (23)
The expectation value of any observable at arbitrary time must be equal across pictures:
$\langle\Psi_{\mathrm{H}}|\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})|\Psi_{\mathrm{H}}\rangle$ $=\langle\Psi_{\mathrm{S}}(t)|\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})|\Psi_{\mathrm{S}}(t)\rangle.$ (24)
Since the quantum state evolves as
$|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{S}}(t,t_{0})|\Psi_{\mathrm{H}}\rangle,$ (25)
the Heisenberg operator is related to its Schrödinger counterpart by
$\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$ $=\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{U}_{\mathrm{S}}(t,t_{0}),$ (26)
which establishes both the connection to the Schrödinger picture and the Heisenberg time evolution.
In the Heisenberg picture, constructing the Heisenberg equation of motion is essential. This requires rewriting the Hamiltonian in Heisenberg picture. Using the relation for $\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$, we define
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0})$
$=\int d^{3}\boldsymbol{x}\hat{D}_{\mathrm{H}}(\alpha,t)\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{H}}(t,\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{H}}(t,\boldsymbol{x}),\hat{A}_{\mathrm{H}}(t,\boldsymbol{x}))\hat{D}_{\mathrm{H}}(\alpha,t)$ (27)
[…]
where, $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]\hat{D}_{\mathrm{S}}(\alpha)$ is employed. In this transformation, the c-number field $\mathcal{A}$ recovers its time-dependence. This Hamiltonian implies that the Dirac field interacts with the sum of the quantized EM field and the coherent laser background: $\hat{A}_{\mathrm{H}}(x)+\mathcal{A}(x)\hat{\mathbb{I}}$. Here the argument $t$ in $\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ labels the Heisenberg-picture operator rather than an explicit time dependence of the Hamiltonian itself, as confirmed by
$\frac{d\hat{H}_{(\mathcal{A})\mathrm{H}}(t)}{dt}$ $=[\partial_{t}\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})]H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0})$
$\quad+\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}][\partial_{t}\hat{U}_{\mathrm{S}}(t,t_{0})]$
$=0.$ (29)
That is, the Hamiltonian
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{H}_{(\mathcal{A})\mathrm{H}}(t_{0})=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] (30)
is time independent. It follows that the Heisenberg picture time-evolution operator is
[…]
consisting only Heisenberg field operators. We can identify $\hat{U}(t,t_{0})$ and $\hat{U}_{\mathrm{S}}(t,t_{0})$: substitute $\hat{O}_{\mathrm{S}}=\hat{U}_{\mathrm{S}}(t,t_{0})$ into Eq. (26). This allows the time evolution operator including the $\mathcal{A}$-effects to be applied consistently in both the Heisenberg and Schrödinger pictures.
4 Furry picture
In this section, we construct the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009] commonly employed in strong-field QED starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified.
One important point should be emphasized at the outset. While the Furry picture provides a powerful and widely used framework for describing a broad class of strong-field QED processes, it is not suitable for capturing effects associated with laser depletion. Such effects lie beyond any formulation based on a fixed background field, including a description in terms of a single coherent state.
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. [Ilderton-Seipt2018], will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
4.1 Basic structures
In the Heisenberg picture, the equation of motion for the field $\hat{\psi}_{\mathrm{H}}$ is
$i\hbar\partial_{t}\hat{\psi}_{\mathrm{H}}(x)$ $=\hat{\psi}_{\mathrm{H}}(x)\hat{H}_{(\mathcal{A})\mathrm{H}}(t)-\hat{H}_{(\mathcal{A})\mathrm{H}}(t)\hat{\psi}_{\mathrm{H}}(x).$ (32)
Although the time variable is written explicitly, from the previous discussion, $\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ is time-independent (see Eq. (30)). We now separate
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)+\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t),$ (33)
where the first term collects the free propagation of the Dirac and EM fields together with the interaction of the Dirac field with the coherent laser background, while the second term represents purely quantum EM interactions. Despite Eq. (30), each of $\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)$ and $\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t)$ carries explicit time dependence.
We now define the field operator transformation as
$\hat{\psi}_{\mathrm{F}}(x)$ $=\hat{V}(t,t_{0})\hat{\psi}_{\mathrm{H}}(x)\hat{V}^{\dagger}(t,t_{0}).$ (34)
Suppose the relation for $\hat{V}(t,t_{0})$ as
$i\hbar\partial_{t}\hat{V}(t,t_{0})$ $=\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t)\hat{V}(t,t_{0})$ (35)
so that the formal solution is
$\hat{V}(t,t_{0})$ $=\mathcal{T}\exp\bigg[-\frac{i}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\,\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t^{\prime})\bigg],$ (36)
with the time-ordering $\mathcal{T}$ and the transformed interaction Hamiltonian
$\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t)$ $=\hat{V}(t,t_{0})\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t)\hat{V}^{\dagger}(t,t_{0}).$ (37)
Under the given definition, the field equation of $\hat{\psi}_{\mathrm{H}}$ becomes the transformed field equation of $\hat{\psi}_{\mathrm{F}}$:
$i\hbar\partial_{t}\hat{\psi}_{\mathrm{F}}(x)$ $=[\hat{\psi}_{\mathrm{F}}(x),\hat{H}_{\mathrm{F}}^{0}(t)]$ (38)
with the transformed free Hamiltonian
$\hat{H}_{\mathrm{F}}^{0}(t)$ $=\hat{V}(t,t_{0})\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)\hat{V}^{\dagger}(t,t_{0}).$ (39)
See also Sect. 4.2 about Eq. (38). We do the same for $\hat{A}_{\mathrm{H}}^{\mu}$, the field equation of $\hat{A}_{\mathrm{F}}^{\mu}$ is imposed.
Consistency of expectation values across pictures requires
$\langle\Psi_{\mathrm{H}}|\hat{\psi}_{\mathrm{H}}(x)|\Psi_{\mathrm{H}}\rangle$ $=\langle\Psi_{\mathrm{F}}(t)|\hat{\psi}_{\mathrm{F}}(x)|\Psi_{\mathrm{F}}(t)\rangle.$ (40)
Thus, the time evolution of the state $|\Psi_{\mathrm{F}}(t)\rangle$ is
$|\Psi_{\mathrm{F}}(t)\rangle$ $=\hat{V}(t,t_{0})|\Psi_{\mathrm{H}}\rangle$
$=\mathcal{T}\exp\bigg[-\frac{i}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t^{\prime})\bigg]|\Psi_{\mathrm{H}}\rangle,$ (41)
[…]
ensuring consistency, with the quantum state and operators matched to the Heisenberg picture at $t_{0}$.
In practical scattering calculations one takes $t\rightarrow\infty$ and $t_{0}\rightarrow$-$\infty$ and the corresponding S-matrix is
$\hat{S}_{\mathrm{F}}$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\hat{V}(t,t_{0})$
[…]
This formalism defines the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering [Ritus1985, Brown-Kibble1964, Nikishov-Ritus1964, Nikishov-Ritus1967] and nonlinear Breit–Wheeler processes [Ritus1985, Nikishov-Ritus1964, Nikishov-Ritus1967, Reiss1962], as confirmed in numerous previous studies.
4.2 Volkov Solutions
We now focus on the Dirac field equation (Eq. (38)) in the Furry picture. Taking e=|e|-e=-|e| and applying the explicit form of the Hamiltonian $\hat{H}_{\mathrm{F}}^{0}(t)$, Eq. (38) becomes
$i\hbar\gamma^{\mu}\bigg[\partial_{\mu}-i\frac{e}{\hbar}\mathcal{A}_{\mu}(x)\bigg]\hat{\psi}_{\mathrm{F}}(x)-mc\hat{\psi}_{\mathrm{F}}(x)$ $=0$ (44)
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions [Joachain_Kylstra_Potvliege_2011, Ritus1985, Piazza2012, Fedotov2023, Volkov1936, Itzykson-Zuber1980]. In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
5 Applications
In this section, we apply the coherent-field QED formalism developed above to two representative phenomena: the Heisenberg–Euler vacuum and laser depletion.
5.1 Heisenberg-Euler vacuum model
The QED vacuum described by the Heisenberg-Euler (HE) effective action is a central concept in strong-field QED and high-intensity laser science. Within the Furry picture, the vacuum-vacuum transition amplitude in a classical background field $\mathcal{A}$ is evaluated as
$Z[\mathcal{A}]=e^{iW[\mathcal{A}]/\hbar}$ $=\langle 0_{\mathrm{F}}|\hat{V}(\infty,\infty)|0_{\mathrm{F}}\rangle,$ (45)
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density [Greiner_Muller_Rafelski_1985, Fedotov2023, Ritus1985, Schwinger1951, Heisenberg-Euler1936, Weisskopf1936, Dunne2005]. Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
$\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle\alpha_{\mathrm{S}}(t_{0})|\alpha_{\mathrm{S}}(t)\rangle$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle 0_{\mathrm{S}}(t_{0})|\hat{U}(t,t_{0})|0_{\mathrm{S}}(t_{0})\rangle$
$=e^{-iC[\mathcal{A}]}\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle.$ (46)
Here, $C[\mathcal{A}]$ denotes the overall c-number phase generated by the action of the background-dressed free Hamiltonian on the vacuum $|0_{\mathrm{F}}\rangle$, i.e., the part of the Hamiltonian that governs vacuum propagation in the presence of the classical background but does not generate particle excitations, analogous to the vacuum energy phase in the Gell-Mann–Low theorem. This phase does not contribute to the Heisenberg–Euler effective action and drops out of all physical observables. A conceptual advantage of the coherent state formulation is that a background field automatically satisfies the Lorenz gauge condition $\partial_{\mu}\mathcal{A}^{\mu}(x)=0$ through the Gupta–Bleuler condition applied to the underlying photon coherent state.
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states.
5.2 Beyond Furry picture
The present method also allows us to describe depletion effects of a coherent laser EM field. In the Furry picture, quantum transitions are treated under a fixed coherent state $\alpha$. Depletion effects corresponds to quantum processes in which a coherent state $\alpha$ evolves into another coherent state $\alpha^{\prime}$. In the Schrödinger picture, the corresponding transition amplitude is
$S_{fi}$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle\alpha_{\mathrm{S}}^{\prime}(t_{0}),\Psi_{\mathrm{S}}^{\prime}(t_{0})|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$
$=\langle\Psi_{\mathrm{S}}^{\prime}(-\infty)|\hat{D}^{\dagger}(\alpha^{\prime})\hat{U}_{\mathrm{S}}(\infty,-\infty)\hat{D}(\alpha)|\Psi_{\mathrm{S}}(-\infty)\rangle.$ (47)
An equivalent form can be found in the work of Ref. [Fedotov2023, Ilderton-Seipt2018]. The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion).
6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories [BRS1976, Tyutin1975], ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition [Gupta1950, Bleuler1950], which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
From the BRST perspective, a physical state $\mid\mathrm{Physical}_{\mathrm{H}}\rangle$ is defined as an element of the cohomology of the nilpotent BRST charge, $\hat{Q}_{(\mathrm{BRST})\mathrm{H}}$, a conserved operator in the Heisenberg picture, satisfying
$\hat{Q}_{(\mathrm{BRST})\mathrm{H}}\mid\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (48)
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field [Kugo-Ojima1979, Nakanishi-Ojima1990]. The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
Importantly, the displacement-based construction employed here acts at the operator level and does not alter the gauge structure of the theory. As a result, the identification of a classical background field with the expectation value of a coherent quantum state remains compatible with the BRST formulation. This observation suggests that the present approach may be naturally extended to gauge theories formulated explicitly in the BRST language, without altering its essential physical content.
7 Path integral representation
In this section, we reformulate the coherent-field QED amplitudes in the path integral representation, showing how the standard background-field generating functional emerges from coherent-state boundary conditions. Throughout this section, $\mathcal{A}$ is treated as a classical field whose status depends on boundary conditions: it is fixed in the Furry picture limit, but becomes a dynamical integration variable when depletion is included. We rewrite the scattering amplitude
$S_{fi}$ $=\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ (49)
introduced in Eq. (47) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture.
The starting point is the observation made in the depletion analysis [Fedotov2023, Ilderton-Seipt2018]: the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed-background limit of coherent-field QED which implicitly assumes a fixed background field and hence $\alpha^{\prime}=\alpha$.
To obtain a path integral representation, recall Eq. (7), we separate the total gauge field into two components,
$\hat{A}_{(\mathrm{tot})\mathrm{H}}^{\mu}(x)$ $=\hat{A}_{\mathrm{H}}^{\mu}(x)+\mathcal{A}^{\mu}(x)\hat{\mathbb{I}}$, (50)
where $\mathcal{A}^{\mu}(x)$ denotes the macroscopic laser field identified with the expectation value of a coherent state, while $\hat{A}_{\mathrm{H}}(x)$ represents the remaining quantum fluctuations of the EM field. The separation is implemented at the level of boundary conditions: the coherent field $\mathcal{A}^{\mu}(x)$ is constrained to interpolate between the classical limits determined by the asymptotic coherent states,
$\mathcal{A}_{(-\infty)}^{\mu}(x)=\lim_{t\to-\infty}\mathcal{A}^{\mu}(x)$ $\sim\langle\alpha_{(\mathrm{in})\mathrm{H}}|\hat{A}_{(\mathrm{in})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{in})\mathrm{H}}\rangle,$ (51)
$\mathcal{A}_{(+\infty)}^{\mu}(x)=\lim_{t\to+\infty}\mathcal{A}^{\mu}(x)$ $\sim\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}|\hat{A}_{(\mathrm{out})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}\rangle,$ (52)
for the asymptotic quantized EM filed $\hat{A}_{(\mathrm{in/out})\mathrm{H}}^{\mu}(x)$. In contrast, the quantied EM field other than the laser field and the Dirac field are treated with standard in–out boundary conditions.
With this separation, the transition amplitude can be written formally as
$\hskip 0.0pt\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\mathcal{D}\bar{\psi}$
$\quad\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (53)
where the functional integration over $\mathcal{A}$ is restricted to field configurations consistent with the coherent state boundary conditions specified above. The Lagrangian density $\mathcal{L}_{\mathrm{QED}}$ is related in the standard way to the Hamiltonian density $\mathcal{H}_{\mathrm{QED}}$ employed in the previous operator-based formulation. This expression provides a first-principles path integral formulation of laser depletion and backreaction, without double counting of photon degrees of freedom.
The conventional generating functional of strong-field QED [Fedotov2023] is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$\hskip 0.0ptZ[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (54)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action [Greiner_Muller_Rafelski_1985, Schwinger1951]. From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler [Gupta1950, Bleuler1950]/BRST [BRS1976, Tyutin1975, Kugo-Ojima1979, Nakanishi-Ojima1990]) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.
8 Conclusion
We have presented a first-principles formulation of QED in which classical background EM fields arise naturally from coherent states of the quantized field. By systematically separating coherent laser modes from other quantum degrees of freedom, our approach unifies operator-based, Furry picture, and path integral descriptions, while maintaining full quantum consistency.
This framework clarifies the conceptual origin of background fields, automatically enforces gauge constraints via the Gupta–Bleuler condition (or equivalently BRST cohomology), and allows treatment of effects beyond the fixed background approximation, including laser depletion and backreaction. Conventional generating functionals and standard strong-field QED formulations appear as well-defined limiting cases within this approach.
The central new elements of this work are the derivation of classical background fields from coherent states at the level of first principles, the construction of QED with coherent background fields within a framework unifying operator-based and path-integral formalisms, and the identification of the fixed background approximation and its extensions, including depletion and backreaction, as natural limiting cases of this construction.
Overall, our results provide a general, intensity-independent foundation for QED in background fields, offering both conceptual clarity and a practical starting point for analyzing high-intensity laser-matter interactions and related strong-field phenomena. In future work, this framework may be extended to study a broader class of quantum processes and to develop systematic approaches for incorporating dynamical background fields in various settings.
Acknowledgements
This work was supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research No. 24K06990 from MEXT of Japan.