### v3

# 322-4 Oroshi-cho, Toki, Gifu, 509-5292, Japan

Abstract
QED formulated in prescribed classical background electromagnetic fields is a standard framework for strong-field and laser–matter interactions. It is usually treated as a theory modified by externally imposed fields, obscuring its precise relation to full QED and, in particular, the role of asymptotic boundary conditions for the gauge field. Starting from an operator-based formulation, we show that QED with background fields is not a distinct theory but arises as a well-defined boundary-condition limit of full QED, in which the classical background field emerges from coherent-state boundary conditions on the quantized electromagnetic field. In this limit, the conventional generating functional used in calculations with prescribed background fields is recovered naturally, while relaxing the boundary conditions allows depletion and backreaction effects to be incorporated within the same framework. The central new result is a rigorous operator-level proof of the equivalence between the fixed background-field approximation and coherent-state asymptotic boundary conditions—a formulation that, to our knowledge, has not been made explicit in previous approaches. We further demonstrate that the apparent time dependence of background fields does not originate from an explicitly time-dependent Hamiltonian, but instead reflects the choice of picture—Schrödinger versus Heisenberg— in the underlying quantum theory. Rather than introducing new properties of coherent states, our analysis provides a first-principles reinterpretation of the fixed background-field approximation as a controlled and picture-dependent limit of full QED.
1 Introduction
Classical background electromagnetic (EM) fields are a central ingredient in quantum electrodynamics (QED), underpinning a wide range of applications from laser–matter interactions to strong-field and external-field problems Saiamin et al. (2006); Joachain et al. (2011); Raicher et al. (2014); Greiner et al. (1985); Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023). In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations of the EM and matter fields are quantized on top of it. In this sense, one works within a formulation of QED in which classical and quantized EM fields coexist. This framework has proven extraordinarily successful in practice, enabling nonperturbative treatments of intense fields and the description of phenomena such as nonlinear Compton scattering Ritus (1985); Brown and Kibble (1964); Nikishov and Ritus (1964, 1967), Breit–Wheeler pair production Ritus (1985); Nikishov and Ritus (1964, 1967); Reiss (1962), and vacuum polarization effects Greiner et al. (1985); Fradkin et al. (1991).
Despite this success, the conceptual status of the fixed background-field approximation remains subtle. In this paper, we use the term “fixed background-field approximation” to denote formulations of QED in which the EM field is decomposed into a prescribed classical background and quantized fluctuations, with the background treated as fixed and unaffected by the quantum dynamics. In conventional formulations—often summarized under the heading of QED in external EM fields or the Furry picture Furry (1951); Moortgat-Pick (2009)—the background field is introduced as an explicit c-number input, frequently endowed with spacetime dependence from the outset. While operationally well defined, this procedure leaves open a fundamental question: how does a classical background EM field arise from the underlying quantum theory of the field, and in what precise sense is the fixed background an approximation?
A natural bridge between fully quantized gauge fields and classical backgrounds is provided by coherent states of the radiation field Glauber (1963b, a); Sudarshan (1963). Coherent states possess classical-like expectation values while remaining genuine quantum states, and have long been recognized as an appropriate description of idealized laser fields Sargent III et al. (1974); Loudon (2000). It is also well known that scattering amplitudes involving coherent states can be reexpressed as amplitudes computed in the presence of background fields, a correspondence discussed in various contexts since early work such as Ref. Frantz (1965) and in later analyses of backreaction and depletion effects Ilderton and Seipt (2018). These observations suggest that the role of background fields in scattering processes may ultimately be tied to the choice of asymptotic boundary conditions for the quantized EM field.
The central claim of this paper is that QED with prescribed classical background fields is not a distinct or modified theory. Rather, the fixed background-field approximation can be identified unambiguously as a controlled boundary-condition limit of full QED. We show that QED with prescribed background fields arises naturally as a limiting case of scattering between coherent states of the quantized EM field, corresponding to fixed coherent-state boundary conditions. In this sense, the background field is not a fundamental input, but an emergent classical object whose status and limitations can be precisely identified.
Our formulation is developed at the operator level, starting from the Schrödinger and Heisenberg pictures of QED Itzykson and Zuber (1980); Sakurai (1967). This perspective makes it possible to disentangle two conceptually distinct issues that are often conflated in standard treatments. First, the classical background field is identified with the expectation value of the quantized gauge field in a coherent state, subject to appropriate physical-state conditions, here implemented via the Gupta–Bleuler formalism Gupta (1950); Bleuler (1950) or, equivalently, BRST cohomology Becchi et al. (1976); Tyutin (1975); Kugo and Ojima (1979). Second, the spacetime dependence of the background field—commonly taken as an external input—is shown to arise from the choice of picture: while the Hamiltonian remains time independent in the Schrödinger picture, the Heisenberg-picture evolution of operators induces a time-dependent background expectation value. In particular, the time dependence of background fields is not introduced ad hoc at the level of the Hamiltonian, but emerges naturally from the underlying quantum dynamics.
Within this framework, the conventional generating functional of QED with a prescribed background field Greiner et al. (1985); Schwinger (1951) is recovered as a special case corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic and fully quantum-mechanical description of effects beyond the fixed background-field approximation, such as depletion and backreaction Ilderton and Seipt (2018), without double counting of photon degrees of freedom. For clarity, we refer to this operator-based formulation—in which classical background fields arise as coherent-state expectation values subject to fixed coherent-state asymptotic boundary conditions—as coherent-field QED. More recently, coherent states have been employed as a technical ingredient to relate amplitudes on background fields to gravitational amplitudes via the classical double copy Ilderton and Lindved (2025). Our aim here is different: rather than exploiting this equivalence as a computational tool, we reinterpret the fixed background-field approximation itself as a limiting procedure based on coherent-state boundary conditions. While operator-based approaches to QED with prescribed background fields—most notably the unstable-vacuum formalism of Fradkin, Gitman and Shvartsman Fradkin et al. (1991), and related analyses by Gavrilov and Gitman Gavrilov and Gitman (1990)—demonstrate that coherent states can mimic prescribed classical fields in specially chosen initial states, these works formulate the background as an externally imposed c-number field and treat the coherent-state displacement operator within the interaction picture. Consequently, the origin of the background field is picture-dependent and not tied to a boundary-condition prescription in full QED. In contrast, the present work identifies the formulation of QED with prescribed background fields itself as a well-defined coherent-state asymptotic boundary-condition limit, formulated at the operator level and independent of picture choice. This allows a first-principles reinterpretation of QED with prescribed background fields and clarifies its precise relation to the underlying quantum theory.
This paper is organized as follows. In Sec. 2 we introduce coherent states of the EM field. Section 3 develops the operator-based formalism connecting the Schrödinger and Heisenberg pictures. In Sec. 4 we discuss the relation to the Furry picture and Volkov solutions, while Sec. 5 presents applications to the Heisenberg–Euler effective action and laser depletion effects. Section 6 comments on the relation to BRST quantization, and Sec. 7 gives the corresponding path-integral formulation based on coherent-state boundary conditions. We conclude in Sec. 8.
2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers to objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition Gupta (1950); Bleuler (1950). Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
$\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}(x)|\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (1)
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adopt the Gupta–Bleuler condition for a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state Glauber (1963b, a); Sudarshan (1963); Sargent III et al. (1974); Loudon (2000) is specified by the following eigen equation
$\hat{A}_{(+)\mathrm{H}}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle$ $=\mathcal{A}_{(+)}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle,$ (2)
and that conjugate equation is $\langle\alpha_{\mathrm{H}}|\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=\langle\alpha_{\mathrm{H}}|\mathcal{A}_{(-)}^{\mu}(x)$ for $\mathcal{A}_{(-)}^{\mu}(x)=[\mathcal{A}_{(+)}^{\mu}(x)]^{*}$ which is the complex conjugate of $\mathcal{A}_{(+)}^{\mu}(x)$. Hence, the expectation value of the EM field in a coherent state is
$\mathcal{A}^{\mu}(x)$ $=\langle\alpha_{\mathrm{H}}|\hat{A}_{\mathrm{H}}^{\mu}(x)|\alpha_{\mathrm{H}}\rangle$ (3)
$=\mathcal{A}_{(+)}^{\mu}(x)+\mathcal{A}_{(-)}^{\mu}(x),$ (4)
which, by identifying $|\mathrm{Physical}_{\mathrm{H}}\rangle=|\alpha_{\mathrm{H}}\rangle$ in the Gupta–Bleuler condition (1), automatically satisfies the Lorenz gauge condition
$\partial_{\mu}\mathcal{A}^{\mu}(x)$ $=0.$ (5)
The coherent state can be generated from the vacuum state $|0_{\mathrm{H}}\rangle$ using the displacement operator $\hat{D}_{\mathrm{H}}(\alpha,t)$, so that
$|\alpha_{\mathrm{H}}\rangle$ $=\hat{D}_{\mathrm{H}}(\alpha,t)|0_{\mathrm{H}}\rangle.$ (6)
Applying the displacement operator to the EM field operator yields
$\hat{D}_{\mathrm{H}}^{\dagger}(\alpha,t)\hat{A}_{\mathrm{H}}^{\mu}(x)\hat{D}_{\mathrm{H}}(t,\alpha)$ $=\hat{A}_{\mathrm{H}}^{\mu}(x)+\mathcal{A}^{\mu}(x)\hat{\mathbb{I}}$, (7)
where, $\hat{\mathbb{I}}$ is the identity operator Glauber (1963a); Loudon (2000); Frantz (1965). In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
3 Basic formalism of coherent-field QED
The purpose of this section is purely conceptual. In standard formulations of QED with prescribed background fields, the time dependence of the background is introduced directly at the level of a c-number function $\mathcal{A}^{\mu}(x)$, obscuring its origin in the underlying quantum theory. Here we show that, once background fields are reinterpreted as expectation values in coherent states, the apparent time dependence of the background field does not originate from a time-dependent Hamiltonian, but exclusively from the choice of picture.
We formulate the incorporation of a classical background field within an operator-based framework, emphasizing the role of the Schrödinger and Heisenberg pictures. We first construct a coherent background field in the Schrödinger picture via a displacement operator acting on the quantized gauge field, which leads to a clean separation between classical and quantum degrees of freedom at the operator level. We then show how the time-dependent background fields familiar from coherent-field QED arise naturally upon transforming to the Heisenberg picture. In this way, the time dependence of the background field is identified as a picture-dependent feature, rather than as an intrinsic property of the Hamiltonian.
Related operator-based approaches to QED with prescribed background fields using coherent-state boundary conditions were discussed previously, notably in Ref. Fradkin et al. (1991); Gavrilov and Gitman (1990). In those formulations, asymptotic coherent states are introduced via a displacement operator described as time “independent,” implicitly within an interaction- or asymptotic-picture setting. Here, we make the role of picture dependence explicit by showing how the time dependence of the background field arises from operator evolution in the Heisenberg picture while remaining absent in the Schrödinger picture by construction.
In strong-field QED/nonlinear QED experiments at high-intensity laser facilities Gales et al. (2018); Mirzaie et al. (2024); Sarri et al. (2025), scattering processes are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background Greiner et al. (1985); Ritus (1985); Di Piazza et al. (2012); Moortgat-Pick (2009); Hartin (2018). It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian. From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009). In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. The detailed construction of the Furry picture in coherent-field QED will be developed in Sect. 4.
3.1 Departure from the Schrödinger picture
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states Itzykson and Zuber (1980); Sakurai (1967). Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
We consider QED in the presence of a coherent background EM field and reformulate it in terms of coherent-field QED. To this end, we assume that a coherent background laser field is included from the outset. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
$|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{D}_{\mathrm{S}}(\alpha)|\Psi_{\mathrm{S}}(t)\rangle,$ (8)
corresponding to a background field description prior to imposing fixed boundary conditions.
The “usual” QED Hamiltonian and its density for the state $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ in the Schrödinger picture are expressed abstractly as
$\hat{H}_{\mathrm{S}}$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$ (9)
$=\int d^{3}\boldsymbol{x}\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{S}}(\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{S}}(\boldsymbol{x}),\hat{A}_{\mathrm{S}}(\boldsymbol{x})),$ (10)
$\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ $=\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{S}}(\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{S}}(\boldsymbol{x}),\hat{A}_{\mathrm{S}}(\boldsymbol{x})).$ (11)
Under this preparation, the Schrödinger equation governing the “coherent laser field plus other particles” state is
$i\hbar\partial_{t}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{H}_{\mathrm{S}}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle,$ (12)
$|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{usual\,QED}}(t,t_{0})|\alpha_{\mathrm{S}}(t_{0}),\Psi_{\mathrm{S}}(t_{0})\rangle,$ (13)
$\hat{U}_{\mathrm{usual\,QED}}(t,t_{0})$ $=\exp\bigg[-\frac{i}{\hbar}\hat{H}_{\mathrm{S}}(t-t_{0})\bigg].$ (14)
We now separate $|\alpha_{\mathrm{S}}(t)\rangle$ and focus on the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$. Substituting Eq. (8) into Eq. (12), the evolution equation becomes
$i\hbar\partial_{t}|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{H}_{\mathrm{S}}\hat{D}_{\mathrm{S}}(\alpha)|\Psi_{\mathrm{S}}(t)\rangle$
$=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}\hat{D}_{\mathrm{S}}(\alpha)]|\Psi_{\mathrm{S}}(t)\rangle.$ (15)
Using the formula (7) in the present picture,
$\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$ $=\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})+\mathcal{A}(t_{0},\boldsymbol{x})\hat{\mathbb{I}}$, (16)
$\mathcal{A}(t_{0},\boldsymbol{x})$ $=\langle\alpha_{\mathrm{S}}(t_{0})|\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})|\alpha_{\mathrm{S}}(t_{0})\rangle,$ (17)
we conclude
$i\hbar\partial_{t}|\Psi_{\mathrm{S}}(t)\rangle$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]|\Psi_{\mathrm{S}}(t)\rangle.$ (18)
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) Gupta (1950); Bleuler (1950) evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture Itzykson and Zuber (1980); Sakurai (1967), the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
$|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{S}}(t,t_{0})|\Psi_{\mathrm{S}}(t_{0})\rangle.$ (19)
The associated evolution operator satisfies
$i\hbar\partial_{t}\hat{U}_{\mathrm{S}}(t,t_{0})$ $=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0}),$ (20)
with the formal solution
$\hat{U}_{\mathrm{S}}(t,t_{0})$ $=\exp\bigg[-\frac{i}{\hbar}H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}](t-t_{0})\bigg].$ (21)
Thus the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$, with the coherent laser field factored out, incorporates the influence of the laser EM field ($\mathcal{A}$) through its explicit appearance in the evolution operator.
This confirms that the Schrödinger picture evolution operator consistently incorporates the coherent background field through displacement, while maintaining full quantum dynamics for all other degrees of freedom. In this way, the framework provides a rigorous starting point for connecting the Schrödinger picture with the Heisenberg picture in the presence of a coherent laser background.
The key point of this subsection is that the usual formulation of QED in prescribed background fields emerges directly from the coherent-field QED framework once the coherent state is factored out in the Schrödinger picture. In particular, starting from the fully quantum description $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$, the displacement by a coherent state produces the familiar fixed background description $|\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] widely used in strong-field QED. This shows that the fixed background-field approximation is not the starting point but rather a derived limit within coherent-field QED. Moreover, because depletion and path-integral formulations require transitions between different coherent states, the fundamental structure of this work is the underlying coherent-field QED structure defined by
$\big(\,|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle,\,\,$ $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]\,\big),$ (22)
not its fixed background reduction.
3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As mentioned earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified Itzykson and Zuber (1980); Sakurai (1967). Thus
$|\Psi_{\mathrm{H}}\rangle$ $=|\Psi_{\mathrm{S}}(t_{0})\rangle,$ (23)
and for any observable $\hat{O}$,
$\hat{O}_{\mathrm{H}}(t_{0},\boldsymbol{x})$ $=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x}).$ (24)
The expectation value of any observable at arbitrary time must be equal across pictures:
$\langle\Psi_{\mathrm{H}}|\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})|\Psi_{\mathrm{H}}\rangle$ $=\langle\Psi_{\mathrm{S}}(t)|\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})|\Psi_{\mathrm{S}}(t)\rangle.$ (25)
Since the quantum state evolves as
$|\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{U}_{\mathrm{S}}(t,t_{0})|\Psi_{\mathrm{H}}\rangle,$ (26)
the Heisenberg operator is related to its Schrödinger counterpart by
$\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$ $=\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{U}_{\mathrm{S}}(t,t_{0}),$ (27)
which establishes both the connection to the Schrödinger picture and the Heisenberg time evolution.
In the Heisenberg picture, constructing the Heisenberg equation of motion is essential. This requires rewriting the Hamiltonian in the Heisenberg picture. Using the relation for $\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$, we define
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0})$
$=\int d^{3}\boldsymbol{x}\hat{D}_{\mathrm{H}}(\alpha,t)\mathcal{H}_{\mathrm{QED}}(\hat{\psi}_{\mathrm{H}}(t,\boldsymbol{x}),\hat{\bar{\psi}}_{\mathrm{H}}(t,\boldsymbol{x}),\hat{A}_{\mathrm{H}}(t,\boldsymbol{x}))\hat{D}_{\mathrm{H}}(\alpha,t)$ (28)
[…]
where, $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]\hat{D}_{\mathrm{S}}(\alpha)$ is employed. In this transformation, the c-number field $\mathcal{A}$ recovers its time-dependence. This Hamiltonian implies that the Dirac field interacts with the sum of the quantized EM field and the coherent laser background: $\hat{A}_{\mathrm{H}}(x)+\mathcal{A}(x)\hat{\mathbb{I}}$. Here the argument $t$ in $\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ labels the Heisenberg-picture operator rather than an explicit time dependence of the Hamiltonian itself, as confirmed by
$\frac{d\hat{H}_{(\mathcal{A})\mathrm{H}}(t)}{dt}$ $=[\partial_{t}\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})]H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}]\hat{U}_{\mathrm{S}}(t,t_{0})$
$\quad+\hat{U}_{\mathrm{S}}^{\dagger}(t,t_{0})H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}][\partial_{t}\hat{U}_{\mathrm{S}}(t,t_{0})]$
$=0.$ (30)
That is, the Hamiltonian
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{H}_{(\mathcal{A})\mathrm{H}}(t_{0})=H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] (31)
is time independent. It follows that the Heisenberg picture time-evolution operator is
[…]
consisting only of Heisenberg field operators. We can identify $\hat{U}(t,t_{0})$ and $\hat{U}_{\mathrm{S}}(t,t_{0})$: substitute $\hat{O}_{\mathrm{S}}=\hat{U}_{\mathrm{S}}(t,t_{0})$ into Eq. (27). This allows the time evolution operator including the $\mathcal{A}$-effects to be applied consistently in both the Heisenberg and Schrödinger pictures.
4 Furry picture
In this section, we construct the Furry picture Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009) commonly employed in strong-field QED (or nonlinear QED) starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified. In particular, while textbook formulation of QED with prescribed external fields postulate a time-dependent c-number background from the outset, here the same time dependence emerges from picture choice rather than being introduced ad hoc at the Hamiltonian level, providing a conceptual link to the standard unstable-vacuum formalism Fradkin et al. (1991).
One important point should be emphasized at the outset. While the Furry picture provides a powerful and widely used framework for describing a broad class of strong-field QED processes, it is not suitable for capturing effects associated with laser depletion. Such effects lie beyond any formulation based on a fixed background field, including a description in terms of a single coherent state.
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. Ilderton and Seipt (2018), will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
4.1 Basic structures
In the Heisenberg picture, the equation of motion for the field $\hat{\psi}_{\mathrm{H}}$ is
$i\hbar\partial_{t}\hat{\psi}_{\mathrm{H}}(x)$ $=\hat{\psi}_{\mathrm{H}}(x)\hat{H}_{(\mathcal{A})\mathrm{H}}(t)-\hat{H}_{(\mathcal{A})\mathrm{H}}(t)\hat{\psi}_{\mathrm{H}}(x).$ (33)
Although the time variable is written explicitly, from the previous discussion, $\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ is time-independent (see Eq. (31)). We now separate
$\hat{H}_{(\mathcal{A})\mathrm{H}}(t)$ $=\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)+\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t),$ (34)
where the first term collects the free propagation of the Dirac and EM fields together with the interaction of the Dirac field with the coherent laser background, while the second term represents purely quantum EM interactions. Despite Eq. (31), each of $\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)$ and $\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t)$ carries explicit time dependence.
We now define the field operator transformation as
$\hat{\psi}_{\mathrm{F}}(x)$ $=\hat{V}(t,t_{0})\hat{\psi}_{\mathrm{H}}(x)\hat{V}^{\dagger}(t,t_{0}).$ (35)
Suppose the relation for $\hat{V}(t,t_{0})$ as
$i\hbar\partial_{t}\hat{V}(t,t_{0})$ $=\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t)\hat{V}(t,t_{0})$ (36)
so that the formal solution is
$\hat{V}(t,t_{0})$ $=\mathcal{T}\exp\bigg[-\frac{i}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\,\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t^{\prime})\bigg],$ (37)
with the time-ordering $\mathcal{T}$ and the transformed interaction Hamiltonian
$\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t)$ $=\hat{V}(t,t_{0})\hat{H}_{(\mathcal{A})\mathrm{H}}^{\mathrm{int}}(t)\hat{V}^{\dagger}(t,t_{0}).$ (38)
Under the given definition, the field equation of $\hat{\psi}_{\mathrm{H}}$ becomes the transformed field equation of $\hat{\psi}_{\mathrm{F}}$:
$i\hbar\partial_{t}\hat{\psi}_{\mathrm{F}}(x)$ $=[\hat{\psi}_{\mathrm{F}}(x),\hat{H}_{\mathrm{F}}^{0}(t)]$ (39)
with the transformed free Hamiltonian
$\hat{H}_{\mathrm{F}}^{0}(t)$ $=\hat{V}(t,t_{0})\hat{H}_{(\mathcal{A})\mathrm{H}}^{0}(t)\hat{V}^{\dagger}(t,t_{0}).$ (40)
See also Sect. 4.2 about Eq. (39). Applying the same transformation for $\hat{A}_{\mathrm{H}}^{\mu}$ yields the field equation of $\hat{A}_{\mathrm{F}}^{\mu}$.
Consistency of expectation values across pictures requires
$\langle\Psi_{\mathrm{H}}|\hat{\psi}_{\mathrm{H}}(x)|\Psi_{\mathrm{H}}\rangle$ $=\langle\Psi_{\mathrm{F}}(t)|\hat{\psi}_{\mathrm{F}}(x)|\Psi_{\mathrm{F}}(t)\rangle.$ (41)
Thus, the time evolution of the state $|\Psi_{\mathrm{F}}(t)\rangle$ is
$|\Psi_{\mathrm{F}}(t)\rangle$ $=\hat{V}(t,t_{0})|\Psi_{\mathrm{H}}\rangle$
$=\mathcal{T}\exp\bigg[-\frac{i}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\hat{H}_{\mathrm{F}}^{\mathrm{int}}(t^{\prime})\bigg]|\Psi_{\mathrm{H}}\rangle,$ (42)
[…]
ensuring consistency, with the quantum state and operators matched to the Heisenberg picture at $t_{0}$.
In practical scattering calculations one takes $t\rightarrow\infty$ and $t_{0}\rightarrow$-$\infty$ and the corresponding S-matrix is
$\hat{S}_{\mathrm{F}}$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\hat{V}(t,t_{0})$
[…]
This formalism defines the Furry picture Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009). Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering Ritus (1985); Brown and Kibble (1964); Nikishov and Ritus (1964, 1967) and nonlinear Breit–Wheeler processes Ritus (1985); Nikishov and Ritus (1964, 1967); Reiss (1962), as confirmed in numerous previous studies.
4.2 Volkov Solutions
We now focus on the Dirac field equation (Eq. (39)) in the Furry picture. Taking the electron charge to be e=|e|-e=-|e| and applying the explicit form of the Hamiltonian $\hat{H}_{\mathrm{F}}^{0}(t)$, Eq. (39) becomes
$i\hbar\gamma^{\mu}\bigg[\partial_{\mu}-i\frac{e}{\hbar}\mathcal{A}_{\mu}(x)\bigg]\hat{\psi}_{\mathrm{F}}(x)-mc\hat{\psi}_{\mathrm{F}}(x)$ $=0.$ (45)
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions Joachain et al. (2011); Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Wolkov (Volkov in English spelling) (1935); Itzykson and Zuber (1980). In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
5 Applications
In this section, we apply the coherent-field QED formalism developed above to two representative phenomena: the Heisenberg–Euler vacuum and laser depletion.
5.1 Heisenberg-Euler vacuum model
The QED vacuum described by the Heisenberg-Euler (HE) effective action is a central concept in strong-field QED and high-intensity laser science. Within the Furry picture, the vacuum-vacuum transition amplitude in a classical background field $\mathcal{A}$ is evaluated as
$Z[\mathcal{A}]=e^{iW[\mathcal{A}]/\hbar}$ $=\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle,$ (46)
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density Greiner et al. (1985); Fedotov et al. (2023); Ritus (1985); Schwinger (1951); Heisenberg and Euler (1936); Weisskopf (1936); Dunne (2005). Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
$\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle\alpha_{\mathrm{S}}(t_{0})|\alpha_{\mathrm{S}}(t)\rangle$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle 0_{\mathrm{S}}(t_{0})|\hat{U}(t,t_{0})|0_{\mathrm{S}}(t_{0})\rangle$
$=e^{-iC[\mathcal{A}]}\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle.$ (47)
Here, $C[\mathcal{A}]$ denotes the overall c-number phase generated by the action of the background-dressed free Hamiltonian on the vacuum $|0_{\mathrm{F}}\rangle$, i.e., the part of the Hamiltonian that governs the vacuum time evolution in the presence of the classical background but does not generate particle excitations, analogous to the vacuum energy phase in the Gell-Mann–Low theorem. This phase does not contribute to the Heisenberg–Euler effective action and drops out of all physical observables. A conceptual advantage of the coherent state formulation is that a background field automatically satisfies the Lorenz gauge condition $\partial_{\mu}\mathcal{A}^{\mu}(x)=0$ through the Gupta–Bleuler condition applied to the underlying photon coherent state.
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states. From the viewpoint of coherent-field QED, the physical interpretation of the HE vacuum becomes more transparent. In the Furry picture the quantity $\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle$ is described as a “vacuum–vacuum transition” in the presence of an externally prescribed background field, a somewhat formal notion in which the vacuum is said to evolve non-trivially despite containing no photons. In contrast, the coherent‑field QED formulation shows that the relevant transition is, more precisely, an $\alpha$-$\alpha$ transition: a coherent photon state representing a laser pulse propagates through the vacuum while retaining its quantum state. Nothing “happens” to the vacuum itself; rather, the effective action arises from the overall phase acquired by a coherent photon state propagating through the vacuum. This interpretation resolves the long-standing conceptual ambiguity of vacuum–vacuum transitions by clarifying that the HE effective action reflects the behavior of such a coherent state, rather than any dynamical evolution of an empty vacuum.
5.2 Beyond Furry picture
The present method also allows us to describe depletion effects of a coherent laser EM field. In the Furry picture, quantum transitions are treated under a fixed coherent state $\alpha$. Depletion effects correspond to quantum processes in which a coherent state $\alpha$ evolves into another coherent state $\alpha^{\prime}$. In the Schrödinger picture, the corresponding transition amplitude is
$S_{fi}$ $=\lim_{t_{0}\rightarrow-\infty,\,t\rightarrow\infty}\langle\alpha_{\mathrm{S}}^{\prime}(t_{0}),\Psi_{\mathrm{S}}^{\prime}(t_{0})|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$
$=\langle\Psi_{\mathrm{S}}^{\prime}(-\infty)|\hat{D}^{\dagger}(\alpha^{\prime})\hat{U}_{\mathrm{S}}(\infty,-\infty)\hat{D}(\alpha)|\Psi_{\mathrm{S}}(-\infty)\rangle.$ (48)
An equivalent form can be found in the work of Ref. Fedotov et al. (2023); Ilderton and Seipt (2018). The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion). We note that depletion modeled via coherent-state deformations has been developed in Ilderton and Seipt (2018). Our treatment complements this by identifying the fixed background-field approximation and its extensions as boundary-condition limits within a single first-principles framework.
6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories Becchi et al. (1976); Tyutin (1975), ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition Gupta (1950); Bleuler (1950), which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
From the BRST perspective, a physical state $\mid\mathrm{Physical}_{\mathrm{H}}\rangle$ is defined as an element of the cohomology of the nilpotent BRST charge, $\hat{Q}_{(\mathrm{BRST})\mathrm{H}}$, a conserved operator in the Heisenberg picture, satisfying
$\hat{Q}_{(\mathrm{BRST})\mathrm{H}}\mid\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (49)
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field Kugo and Ojima (1979); Nakanishi and Ojima (1990). The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
Importantly, the displacement-based construction employed here acts at the operator level and does not alter the gauge structure of the theory. As a result, the identification of a classical background field with the expectation value of a coherent quantum state remains compatible with the BRST formulation. This observation suggests that the present approach may be naturally extended to gauge theories formulated explicitly in the BRST language, without altering its essential physical content.
7 Coherent-state boundary conditions in path-integral representation
In contrast to conventional treatments of QED in prescribed background fields—where a classical field $\mathcal{A}$ is inserted directly into the action—our formulation shows that the same background-field functional integral arises as a boundary-condition limit of the full quantum theory, obtained by imposing coherent-state asymptotic conditions on the gauge field. This identifies the standard background-field path integral as a controlled limit of coherent-field QED rather than as an independent modification of the theory. The following derivation is standard except for the implementation of coherent-state boundary conditions, which encode the classical background field as a limiting case.
We reformulate the coherent-field QED amplitudes in the path-integral representation, showing explicitly how the standard generating functional of strong-field QED emerges from coherent-state boundary conditions. Throughout this section, $\mathcal{A}$ is treated as a classical field whose status depends on boundary conditions: it is fixed in the Furry picture limit, but becomes a dynamical integration variable when depletion is included. We rewrite the scattering amplitude
$S_{fi}$ $=\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ (50)
introduced in Eq. (48) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path-integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture. From the path-integral viewpoint, fixing coherent-state boundary conditions for the EM field yields the conventional generating functional of QED in a prescribed background field as a controlled limit, while allowing $\alpha\rightarrow\alpha^{\prime}$ transitions promotes the background to a dynamical variable integrated over histories consistent with the asymptotic coherent states. This makes explicit how the standard generating functional of strong-field QED and the Heisenberg–Euler effective action arise within our framework, without double counting of photon degrees of freedom Greiner et al. (1985); Schwinger (1951); Fradkin et al. (1991); Fedotov et al. (2023).
The starting point is the observation made in the depletion analysis Fedotov et al. (2023); Ilderton and Seipt (2018): the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed background-field limit of coherent-field QED which implicitly assumes a fixed background-field and hence $\alpha^{\prime}=\alpha$.
To obtain a path-integral representation, recall Eq. (7), we separate the total gauge field into two components,
$\hat{A}_{(\mathrm{tot})\mathrm{H}}^{\mu}(x)$ $=\hat{A}_{\mathrm{H}}^{\mu}(x)+\mathcal{A}^{\mu}(x)\hat{\mathbb{I}}$, (51)
where $\mathcal{A}^{\mu}(x)$ denotes the macroscopic laser field identified with the expectation value of a coherent state, while $\hat{A}_{\mathrm{H}}(x)$ represents the remaining quantum fluctuations of the EM field. The separation is implemented at the level of boundary conditions: the coherent field $\mathcal{A}^{\mu}(x)$ is constrained to interpolate between the classical limits determined by the asymptotic coherent states,
$\mathcal{A}_{(-\infty)}^{\mu}(x)$ $=\langle\alpha_{(\mathrm{in})\mathrm{H}}|\hat{A}_{(\mathrm{in})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{in})\mathrm{H}}\rangle\quad\mathrm{for}\quad t\rightarrow-\infty,$ (52)
$\mathcal{A}_{(+\infty)}^{\mu}(x)$ $=\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}|\hat{A}_{(\mathrm{out})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}\rangle\quad\mathrm{for}\quad t\rightarrow+\infty,$ (53)
for the asymptotic quantized EM field $\hat{A}_{(\mathrm{in/out})\mathrm{H}}^{\mu}(x)$. In contrast, the quantized EM field other than the laser field and the Dirac field are treated with standard in–out boundary conditions.
With this separation, the transition amplitude can be written formally as
$\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}$
$\quad\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (54)
where the functional integration over $\mathcal{A}$ is restricted to field configurations consistent with the coherent state boundary conditions specified above. The Lagrangian density $\mathcal{L}_{\mathrm{QED}}$ is related in the standard way to the Hamiltonian density $\mathcal{H}_{\mathrm{QED}}$ employed in the previous operator-based formulation. This expression provides a systematic operator-based path integral formulation of laser depletion and backreaction, derived directly from operator-level matrix elements, without double counting of photon degrees of freedom.
The conventional generating functional of strong-field QED Fedotov et al. (2023) is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$Z[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (55)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action Greiner et al. (1985); Schwinger (1951). From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path-integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler Gupta (1950); Bleuler (1950)/BRST Becchi et al. (1976); Tyutin (1975); Kugo and Ojima (1979); Nakanishi and Ojima (1990)) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.
8 Conclusion
The central result of this work is that QED with prescribed classical background EM fields is not a distinct or modified theory. Rather, it is a boundary-condition limit of full QED, obtained by fixing coherent-state asymptotic boundary conditions of the EM field. This perspective differs from earlier uses of coherent states, in which the background field was treated as an externally prescribed c-number input rather than as a boundary‑condition limit of full QED.
Starting from an operator formulation of QED, we have shown how the fixed background-field approximation arises naturally as a limiting case of scattering between coherent states, thereby providing a precise and operator-based characterization of its conceptual status within the underlying quantum theory.
Our formulation makes explicit how classical background fields emerge as expectation values of the quantized gauge field in coherent states, while fully preserving the quantum dynamics of the remaining degrees of freedom. Gauge constraints are implemented consistently at the level of the quantized EM field through the Gupta–Bleuler condition, or equivalently through BRST cohomology. Consequently, the coherent-state expectation value defining the classical background field automatically satisfies the corresponding Lorenz-gauge condition. Once identified as a c-number classical field, the background field may therefore be treated in a convenient gauge, without imposing additional physical-state constraints beyond those already enforced in the underlying quantum theory. The time dependence of background fields is shown to arise from the choice of picture, rather than being introduced ad hoc at the level of the Hamiltonian.
Conventional generating functionals and standard formulations of QED with prescribed background fields are recovered as well-defined limiting cases corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic extension beyond the fixed background-field approximation, enabling a consistent quantum-mechanical treatment of effects such as depletion and backreaction without double counting of photon degrees of freedom.
The results presented here do not introduce new properties of coherent states themselves, but instead formulate the conventional use of background fields in QED as a controlled and conceptually transparent approximation. The scope of this work is conceptual rather than computational, and we make no claim regarding improvements to perturbative techniques or explicit amplitudes beyond clarifying the operator-level structure underlying QED with prescribed background fields. At the conceptual level, the coherent-state transitions underlying the fixed background-field approximation are intrinsically nonperturbative in nature. In this sense, the fixed background description of QED appears as a genuinely nonperturbative boundary-condition limit of the full theory. The coherent-state specification of the asymptotic EM field conditions encodes vacuum persistence, depletion and pair creation in a way that does not rely on any perturbative expansion, but instead on the global structure of the Hilbert space.
This perspective provides a solid foundation for the analysis of strong-field and laser–matter interactions within QED, and for future extensions in which background fields are treated dynamically within a fully quantum framework.
Acknowledgements.
This work was supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research No. 24K06990 from MEXT of Japan.